OpenStudy (anonymous):

find the limit of x^2-4/x+2 as x approaches - 2

OpenStudy (anonymous):

factor here

OpenStudy (anonymous):

you have a difference of 2 squares

OpenStudy (anonymous):

x^2-4

OpenStudy (anonymous):

x^2-4=(x+2)(x-2)

OpenStudy (anonymous):

(x+2)(x-2)/(x+2)=x-2

OpenStudy (anonymous):

as x approaches -2, therefore, y approaches -4